You are an engineer Disney World. You expect one of your brand new major attractions will begin having significant maintenance expenses starting at the end of year 4 with a $60,000 expense. This maintenance expense will increase by $4,000 each year after this inital year 4 expense through the end of year 8. How much money would you have to invest today into an account paying 3% interest compounded annually to completely cover all future maintenance costs for this attraction?
Answer #1

A B 1 Year Cf 2 1 $ 3 2 $ 4 3 $ 5 4 $ 60,000.00 6 5 $ 64,000.00 7 6 $ 68,000.00 8 7 $ 72,000.00 9 8 $ 76,000.00 10 Invest tod

Working:-

A B Cf 1 Year 2 1 0 3 2 0 4 3 0 5 4 60000 6 5 =B5+4000 76 =B6+4000 8 7 =B7+4000 98 =B8+4000 10 Invest today =NPV(3%,B2:39)